OpenStudy (anonymous):

Let f(x) = 5x – 26 and g(x) = 6x. Part 1 [4 points] What is (f ◦ g)(x) ? Part 2 [2 points] Use complete sentences to describe the method you used to solve this problem.

OpenStudy (anonymous):

f(g(x)) = f(6x) = 5(6x) - 26 = 30x - 26 f(g(x)) is f of g(x). Thus, we substitute g(x) into the x-variable of f. Next, we expand out the entire term, and then we are done

OpenStudy (waleed_imtiaz):

F[g(x)]=F(6x)=5(6x)-26=30x-26 answer......

OpenStudy (anonymous):

(f ○ g)(x) = f( g(x) ) = f( 6x ) = 5(6x) - 26 = 30x - 26 Given a function of a function, sub in the entire inner function (g(x)) where ever you see x in the outer function (f(x)).

OpenStudy (anonymous):

What would part 2 be exactly

OpenStudy (waleed_imtiaz):

It is defined to be the composition of f of g of x.... written as (f ○ g)(x).